A gas has a volume of 6.85 L at a pressure of 0.650 atm. What is the volume of the gas if the pressure is decreased to 0.435 atm?
1
Expert's answer
2017-04-20T06:58:27-0400
Solution:
The ideal gas is described by the law
PV=nRT,
where P is the pressure, V is the volume, n is the amount in moles, T is the temperature, and R is the ideal gas constant. From this law, the volume, V is
V=nRT/P.
With fixed values of n and T (R is fixed as it is a constant), the volume changes inversely to the pressure:
V_1/V_2 =P_2/P_1 →V_2=P_1/P_2 V_1
Thus,
V_2=P_1/P_2 V_1=0.650/0.435∙6.85L≈10.236L
Answer:
The volume of the gas would be 10.236 L.